how manyu liters of punch containing 50% friut juice must be added to 20 liters of punch containing 20% fruit juice to produce a punch that it #)% fruit juice.
Answer by mananth(16949)

------ Percent ---------------- quantity
Punch 20 ----------------20
Punch 50 ----------------x
Total 30 ----------------20+x
20*20+50*x=30(20+x)
400+50x =600+30 x
50x-30x =600-400
20x=200
/ 20
x=10 liters 50% Punch